OpenStudy (anonymous):

please help to prove cosh 3x = 4 cosh^3 x - 3 cosh x

OpenStudy (vishal_kothari):

do u know the proof for cos3x=4cos^3 x - 3 cos x?

OpenStudy (vishal_kothari):

cosh(a + b) = coshacoshb - sinhasinhb sinh(a + b) = sinhacoshb + coshasinhb sinh^2a + cosh^2a = 1 Consider the left hand side: cosh3x = cosh(2x + x) = cosh2xcoshx - sinh2xsinhx = cosh(x + x)coshx - sinh(x + x)sinhx = (coshxcoshx - sinhxsinhx)coshx - (sinhxcoshx + coshxsinhx)sinhx = (cosh^2x - sinh^2x)coshx - (2sinhxcoshx)sinhx = (cosh^2x - (1 - cosh^2x))coshx - 2sinh^2xcoshx = (cosh^2x - 1 + cosh^2x)coshx - 2(1 - cosh^2x)coshx = (2cosh^2x - 1)coshx - 2(coshx - cosh^3x) = 2cosh^3x - coshx - 2coshx + 2cosh^3x = 4cosh^3x - 3coshx

OpenStudy (vishal_kothari):

i think it will be fine....

OpenStudy (anonymous):

thanks :D this really helpful

OpenStudy (vishal_kothari):

please be my fan...